net force and race cars

Rank the following sports cars by the net force from greatest to least required to produce the given motion. Assume a constant acceleration for each car.
A Lamborghini Gallardo of mass mL = 1430 kg starts with a speed of 9.16 m/s and goes a distance of d = 102 m in a time of t = 4.25 s.
A Bugatti Veyron of mass mB = 1890 kg goes from 15.5 to 65.5 mph over a distance of d = 40.5 m.
A Dodge Viper of mass mD = 1570 kg goes from 15.5 to 85.5 mph in a time of t = 4.08 s.
A Nissan GT-R of mass mN = 1680 kg goes from 11.2 to 38.8 m/s over a distance of d = 68.8 m.
Answer #1

A Lamborghini Gallardo of mass mL = 1430 kg starts with a speed of 9.16 m/s and goes a distance of d = 102 m in a time of t = 4.25 s.

F = m(v^2-u^2)/2S

F= 588N


A Bugatti Veyron of mass mB = 1890 kg goes from 15.5 to 65.5 mph over a distance of d = 40.5 m.


F=19023 N


A Dodge Viper of mass mD = 1570 kg goes from 15.5 to 85.5 mph in a time of t = 4.08 s.


F = 12040 N



A Nissan GT-R of mass mN = 1680 kg goes from 11.2 to 38.8 m/s over a distance of d = 68.8 m.


F = 16849 N


Highest to lowest 


Bugatti Veyron


Nissan GT-R


Dodge Viper 


Lamborghini Gallard0


answered by: Fosauri